# GIST: Gauge-Invariant Spectral Transformers for Scalable Graph Neural Operators

### 17 Mar 2026 v1

@@ 3.1 Preliminaries
GIST computational complexity is dominated by two components. First, spectral embedding computation via FastRP scales as $\mathcal{O}(N\cdot r\cdot k)$ where $r$ is the embedding dimension and $k$ is the number of power iterations. Second, linear transformer blocks with Gauge-Invariant Spectral Self-Attention on $d$-dimensional node features scale as $\mathcal{O}(N\cdot d^{2})$. Overall, this gives end-to-end $\mathcal{O}(N\cdot d^{2}+N\cdot r\cdot k)$ scaling, i.e. linear in the number of nodes $N$. This contrasts with $\mathcal{O}(N^{3})$ for exact eigendecomposition and $\mathcal{O}(N^{2}d)$ for standard quadratic attention. Empirical validation of this linear scaling is provided in Appendix A.4: Figure 3 demonstrates that both VRAM consumption and forward pass time grow linearly with graph size up to 500K nodes on DrivAerNet samples, confirming our theoretical analysis and enabling the large-scale neural operator experiments in Section 4.
@@ 4.1 Theoretical Guarantees: GIST as Discretization-Invariant Neural Operator
We now formalize how GIST’s gauge-invariant design enables discretization-invariant learning with bounded error, a property essential for neural operator applications. Physical problems (computational fluid dynamics, structural mechanics, shape analysis) are defined on continuous manifolds but discretized into computational meshes corresponding to graphs. Different mesh resolutions produce different graph Laplacians with different spectral decompositions, each involving arbitrary gauge choices (sign flips, eigenspace rotations, solver-dependent orderings). Without gauge invariance, parameters trained on one discretization fail to transfer to another, preventing convergence to a continuum limit. Gauge invariance ensures that attention weights converge to a well-defined continuum kernel, enabling provably bounded discretization mismatch error that vanishes as mesh resolution increases.
@@
Outputs of GIST applied to different discretizations of the same $m$-dimensional manifold, obtained by random sampling nodes, converge to each other with error $\mathcal{O}(n^{-1/(m+4)})$ where $n$ is the coarser resolution. This ensures learned parameters transfer across arbitrary mesh resolutions with bounded error that vanishes as resolution increases.
Proof sketch. The result follows from: (1) spectral convergence of graph Laplacians to manifold eigenfunctions (BelkinNiyogi2008; CalderGarciaTrillos2020); (2) random projections preserve inner products (Johnson-Lindenstrauss); (3) gauge-invariant attention depends only on these inner products, ensuring parameter transfer. Full formulation and proof are found in Appendix A. ∎
@@ 4.2 Discretization Invariance Empirical Verification
Table 1: Discretization invariance verification. Models trained on 50% decimated mesh and tested on full-resolution mesh of the same car geometry. Gauge-invariant architecture transfers successfully; non-gauge-invariant baseline fails (we show R-squared, averaged over 6 random seeds, confidence intervals are std. error).
Architecture Train $R^{2}$ $\uparrow$ Test $R^{2}$ $\uparrow$
(coarse mesh) (fine mesh)
[…]
@@ 4.3 Node Classification Tasks
To demonstrate the key advantages of GIST, we evaluate it on both transductive and inductive node classification datasets. Transductive tasks are a common graph neural networks paradigm and consist of training and evaluating the model on the same graph, with the goal of predicting at test time node labels that were not provided at training (infilling). Inductive tasks on the other hand, operate on a disjoint set of graphs and aim to predict properties on an entirely new graph.
@@
We evaluate our method on transductive graph benchmarks using the official training, validation, and test splits and evaluation protocols. For each method, we select optimal hyperparameters by optimizing over the validation split of each dataset.
GIST-specific parameters (FastRP $k$ and $r$) are optimized through HPO. Appendix A.3 shows final accuracy is robust to variation of these parameters, with performance saturating at relatively low $r$ as predicted by Johnson-Lindenstrauss.
For the final result, we train on the combined training and validation set and evaluate the model on the corresponding test set. We train across multiple random seeds and report the mean ± standard deviation of the relevant metric.
@@ 4.3.1 Transductive Tasks
Across these benchmarks, GIST is competitive with strong graph convolutional and transformer-style baselines (see Table 2). On Pubmed, GIST attains the best mean accuracy among the reported methods ($81.20\%\pm 0.41$), narrowly surpassing enhanced GCN variants (e.g., $81.12\%\pm 0.52$) and outperforming GAT/GraphSAGE families. On Cora and Citeseer, GIST achieves results comparable to the top results (within $\sim$1–2 points of GCNII/SGFormer and the enhanced GCN), landing at $84.00\%\pm 0.60$ and $71.31\%\pm 0.50$, respectively.
[…]
@@
[…]
Table 3: Inductive node classification on PPI, Elliptic, Arxiv, and Photo. Results are reported as micro-F1 (higher is better). Benchmark results are taken from the following references (a – means no reported results): (Chen2020; weber2019elliptic; Chiang2019; Velickovic2018; Zhang2018a; Zeng2020; Chen2025; Brody2022; bo2023specformer; rampasek2022GPS)
@@ 4.3.2 Inductive Tasks
We evaluate our method on four inductive benchmarks: PPI, Elliptic, Arxiv, and Photo. PPI is a collection of 24 disjoint tissue-specific protein–protein interaction graphs where nodes (proteins) have 50 features and 121 non–mutually-exclusive GO labels. Following the standard split (20 graphs for training, 2 for validation, and 2 for testing), we report micro-averaged F1 on the unseen test graphs. Elliptic is a time-evolving directed Bitcoin transaction graph with 203,769 transactions (nodes), 234,355 payment-flow edges, and 166 features across 49 snapshots, labeled licit/illicit with many nodes unlabeled due to class imbalance. We train on the first 29 time steps, validate on the next 5, and test on the last 14, reporting micro-F1. For the Arxiv citation graph (269,343 nodes, 1,166,243 edges, 128 features, and 40 classes) and the Amazon Photo co-purchase network (7,650 nodes, 119,081 edges, 745 features, and 8 classes), we evaluate GIST in the inductive node classification setting using official splits (Fey/Lenssen/2019; Fey/etal/2025).
GIST achieves strong performance across all datasets. On PPI, GIST reaches $99.50\%\pm 0.03$ micro-F1, matching the best large-scale sampling methods and deep residual GCNs (see Table 3, on par with GCNIII and within noise of the strongest GCNII setting ($99.53\%$)). On the temporally inductive Elliptic dataset, GIST attains $94.70\%\pm 0.03$ micro-F1. While this trails the strongest GraphSAGE configuration, GIST maintains stable performance across future time steps. On Arxiv, GIST achieves a mean micro-F1 of $72.12\%\pm 0.21$, competitive with recent spectral transformers (PolyFormer: 72.42%, SpecFormer: 72.37%, Exphormer: 72.44%) while maintaining efficient $\mathcal{O}(N)$ scaling through its linear attention formulation. On the smaller but feature-rich Photo graph, GIST attains $94.42\%\pm 0.40$ micro-F1, which is also competitive with recent spectral and polynomial Transformer variants. Note that GraphGPS, goes OOM on ArXiv primarily due to poor scalability properties and the lack of the gauge invariance trick. These findings demonstrate GIST’s effectiveness as a competitive graph learning approach, validating the successful trade-off between computational overhead and representational power.
@@ 4.4 Neural Operators
We now validate GIST’s discretization-invariant properties (Section 4.1) on large-scale mesh-based regression through the use of the DrivAerNet and DrivAerNet++ Dataset (elrefaie2024drivaernet). DrivAerNet is a high-fidelity CFD dataset of parametric car geometries comprising 4,000 designs with approximately 500k surface vertices per car and accompanying aerodynamic fields. This is extended by DrivAerNet++, which increases the scale to over 10,000 designs and introduces greater geometric diversity across multiple vehicle classes, including SUVs, sedans, and hatchbacks. We model each car as a graph whose nodes are surface vertices and edges follow mesh connectivity. Our task is node-level regression of the surface pressure field on previously unseen cars following the published train/validation/test split.
Table 4 shows GIST achieves state-of-the-art on both benchmarks: 20.10% relative $\ell_{2}$ error on DrivAerNet (vs. 20.35% for the previous best) and 18.60% on DrivAerNet++ (vs. 20.05%). GIST’s linear scaling enables direct processing of these 500K-node meshes without downsampling, while prior methods require lossy projection to regular grids or lower-dimensional latent spaces. GIST’s spectral attention provides global receptive fields in a single layer, which aids performance on this task.
@@ 5 Conclusions
We presented GIST, a gauge-invariant spectral transformer that achieves three properties: gauge invariance for principled generalization across graphs and meshes, $\mathcal{O}(N)$ complexity for scalability, and global spectral attention for capturing long-range dependencies. Crucially, while random projections break gauge symmetry in individual embeddings, inner products between projected embeddings remain approximately invariant. By restricting attention to these gauge-invariant inner products, GIST recovers the symmetry algorithmically while maintaining end-to-end linear scaling.
Gauge invariance is essential for neural operator applications: it ensures learned parameters transfer across mesh resolutions with discretization mismatch error $\mathcal{O}(n^{-1/(m+4)})$ that vanishes as resolution increases. Prior spectral methods address gauge invariance or efficiency separately. The novelty of the GIST architecture that we propose is the careful design that successfully combines these properties.
@@
Appendix A Proofs and Technical Details
@@ A.1 Proof of Proposition Proposition: Gauge-Invariant Spectral Self-Attention
We prove that the Gauge-Invariant Spectral Self-Attention mechanism is discretization-invariant with quantifiable discretization mismatch error through three stages of analysis. The proof establishes that the positional encodings used in the attention mechanism converge to the continuum Green’s function, allowing us to bound the discretization mismatch error between any two discretizations of the same manifold.
@@
Gauge-Invariant Spectral Self-Attention is a discretization-invariant Neural Operator with bounded discretization mismatch error. Let $M$ be a compact $m$-dimensional Riemannian manifold and $\mathcal{G}_{n}$ a sequence of graphs obtained by sampling $n$ nodes from $M$ with $n\to\infty$. Let $\phi_{i}^{n}$ be the Laplacian eigenmaps of $\mathcal{G}_{n}$ (equation 2). Then:
@@
    The inner products $\langle\phi_{i}^{n},\phi_{j}^{n}\rangle$ converge to the Green’s function $G_{M}(x_{i},x_{j})$ of the Laplace-Beltrami operator at rate $\mathcal{O}(n^{-1/(m+4)})$.
@@
    For any two discretizations $\mathcal{G}_{n}$ and $\mathcal{G}_{n^{\prime}}$ of the same manifold $M$ with $n\leq n^{\prime}$, the attention kernel values $\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle$ at corresponding points $x_{i},x_{j}\in M$ satisfy
    $\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|=\mathcal{O}(n^{-1/(m+4)}),$
    where $\tilde{\phi}_{i}^{n}=R^{n}\phi_{i}^{n}$ denotes the projected embeddings. This bounded kernel mismatch ensures learned parameters transfer across discretizations (discretization-invariance).

### 10 Apr 2026 v2

@@ 3.1 Preliminaries
GIST computational complexity is dominated by two components. First, spectral embedding computation scales as $\mathcal{O}(N\cdot r\cdot k)$ where $r$ is the embedding dimension and $k$ is the number of power iterations (e.g., $k$ polynomial filter steps for FastRP). Second, linear transformer blocks with Gauge-Invariant Spectral Self-Attention on $d$-dimensional node features scale as $\mathcal{O}(N\cdot d^{2})$. Overall, this gives end-to-end $\mathcal{O}(N\cdot d^{2}+N\cdot r\cdot k)$ scaling, i.e. linear in the number of nodes $N$. This contrasts with $\mathcal{O}(N^{3})$ for exact eigendecomposition and $\mathcal{O}(N^{2}d)$ for standard quadratic attention. This linear scaling is empirically validated in Appendix C.2: Figure 3 demonstrates that both VRAM consumption and forward pass time grow linearly with graph size up to 500K nodes on DrivAerNet samples, confirming our theoretical analysis and enabling the large-scale neural operator experiments in Section 4.
@@ 4.1 Theoretical Guarantees: GIST as Discretization-Invariant Neural Operator
We now formalize how gauge invariance enables discretization-invariant learning with bounded error. Physical problems (computational fluid dynamics, structural mechanics, shape analysis) are discretized into computational meshes (i.e. graphs), with different resolutions producing different graph Laplacians and spectral decompositions involving arbitrary gauge choices (sign flips, eigenspace rotations, solver artifacts). Without gauge invariance, parameters fail to transfer across discretizations; gauge invariance ensures attention weights converge to a well-defined continuum kernel with bounded discretization mismatch error.
@@
Outputs of GIST applied to different discretizations of the same $m$-dimensional manifold, obtained by random sampling nodes, converge to each other with error $\mathcal{O}(n^{-1/(m+4)})+\mathcal{O}(r^{-1/2})$, where $n$ is the coarser resolution and $r$ is the spectral embedding dimension. This ensures learned parameters transfer across arbitrary mesh resolutions with bounded error that vanishes as resolution and embedding dimension increase.
Proof sketch. The result follows from: (1) spectral convergence of graph Laplacians to manifold eigenfunctions at rate $\mathcal{O}(n^{-1/(m+4)})$ (BelkinNiyogi2008; CalderGarciaTrillos2020); (2) spectral embeddings $\widetilde{\Phi}=f(P)\cdot R$ produce inner products that estimate the exactly gauge-invariant kernel $K=f(P)f(P)^{\top}$ with statistical error $\mathcal{O}(r^{-1/2})$ (Lemma A.4, Proposition A.5, Lemma A.6); (3) gauge-invariant attention depends only on these inner products, ensuring parameter transfer with total error $\mathcal{O}(n^{-1/(m+4)})+\mathcal{O}(r^{-1/2})$. Full formulation and proof are found in Appendix A.
@@ 4.2 Discretization Invariance Empirical Verification
Table 1: Discretization invariance verification. Models trained on 50% decimated mesh and tested on full-resolution mesh of the same car geometry. Gauge-invariant architecture transfers successfully; non-gauge-invariant baseline fails (we show R-squared, averaged over 6 random seeds $\pm$ standard deviation).
Architecture Train $R^{2}$ $\uparrow$ Test $R^{2}$ $\uparrow$
(coarse mesh) (fine mesh)
[…]
@@ 4.3 Node Classification Tasks
To demonstrate the key advantages of GIST, we evaluate it on both transductive and inductive node classification datasets. Transductive tasks are a common graph neural networks paradigm and consist of training and evaluating the model on the same graph, with the goal of predicting at test time node labels that were not provided at training (infilling). Inductive tasks instead operate on a disjoint set of graphs and aim to predict properties on a new graph.
@@
We evaluate our method on transductive graph benchmarks using the official training, validation, and test splits and evaluation protocols. For each method, we select optimal hyperparameters by optimizing over the validation split of each dataset. For final results we train on combined training and validation set.
Spectral embedding parameters based on FastRP ($k$ and $r$) are optimized through HPO. Appendix C.1 shows final accuracy is robust to variation, with performance saturating at relatively low $r$, consistent with the $\mathcal{O}(r^{-1/2})$ concentration rate of the kernel estimator (Lemma A.6).
@@ 4.3.1 Transductive Tasks
Table 2: Transductive node classification on the Planetoid benchmarks (Cora, Citeseer, Pubmed). We report test accuracy (%) as average $\pm$ standard deviation across random seeds using the standard public split (higher is better). Benchmark results are taken from the following references (a ‘–’ means no reported results): a(Luo2024), b(Choi2022), c(Chen2020) .
@@
Although GIST’s gauge invariance is most impactful in inductive and large-scale settings (where spectral gauge changes occur between training and testing), it remains competitive on transductive benchmarks, demonstrating the generality of the architecture. Across these benchmarks, GIST is competitive with strong graph convolutional and transformer-style baselines (see Table 2). On Pubmed, GIST attains the best mean accuracy among the reported methods ($81.20\%\pm 0.41$), narrowly surpassing enhanced GCN variants (e.g., $81.12\%\pm 0.52$) and outperforming GAT/GraphSAGE families. On Cora and Citeseer, GIST achieves results comparable to the top results (within $\sim$1–2 points of GCNII/SGFormer and the enhanced GCN), landing at $84.00\%\pm 0.60$ and $71.31\%\pm 0.50$, respectively.
Table 3: Inductive node classification on PPI, Arxiv, and Photo. Results are reported as micro-F1 (higher is better) as average $\pm$ standard deviation across random seeds. Benchmark results are taken from the following references (a ‘–’ means no reported results): a(Zeng2020), b(Brody2022), c(bo2023specformer), d(Velickovic2018), e(Chen2020), f(Choi2022), g(Chiang2019), h(Chen2025), i(shirzad2023exphormer), j(Wu2024c), k(Ma2024), l(Zhang2018a). source paper reports 95% confidence interval as uncertainty interval which is rescaled to standard deviation. source paper does not specify what the reported uncertainty intervals are.
@@ 4.3.2 Inductive Tasks
We evaluate our method on three inductive benchmarks: PPI, Arxiv, and Photo. PPI is a collection of 24 disjoint tissue-specific protein–protein interaction graphs where nodes (proteins) have 50 features and 121 non–mutually-exclusive GO labels. Following the standard split (20 graphs for training, 2 for validation, and 2 for testing), we report micro-averaged F1 on the unseen test graphs. For the Arxiv citation graph (269,343 nodes, 1,166,243 edges, 128 features, and 40 classes) and the Amazon Photo co-purchase network (7,650 nodes, 119,081 edges, 745 features, and 8 classes), we evaluate GIST in the inductive node classification setting using official splits (Fey/Lenssen/2019; Fey/etal/2025).
GIST achieves strong performance across all datasets. On PPI, GIST reaches $99.50\%\pm 0.03$ micro-F1, matching the best large-scale sampling methods and deep residual GCNs (see Table 3, on par with GCNIII and within noise of the strongest GCNII setting ($99.53\%$)). On Arxiv, GIST achieves a mean micro-F1 of $72.12\%\pm 0.21$, competitive with recent spectral transformers (PolyFormer: 72.42%, SpecFormer: 72.37%, Exphormer: 72.44%) while maintaining efficient $\mathcal{O}(N)$ scaling through its linear attention formulation. On the smaller but feature-rich Photo dataset, GIST attains $95.61\pm 0.23$ micro-F1, which slightly surpasses SpecFormer’s state-of-the-art performance of $95.48\pm 0.16$. Note that GraphGPS goes OOM on ArXiv due to poor scalability. These findings demonstrate GIST’s effectiveness as a competitive graph learning approach, validating the successful trade-off between computational overhead and representational power.
@@ 4.4 Neural Operators
We validate GIST’s discretization-invariant properties (Section 4.1) on four surface pressure prediction benchmarks of increasing scale, all involving node-level regression of aerodynamic fields on unseen geometries.
AirfRANS (bonnet2022airfrans) is a high-fidelity computational fluid dynamics dataset of 2D airfoils under Reynolds-Averaged Navier-Stokes (RANS) simulation, comprising 1000 airfoil designs with approximately 200K mesh nodes per sample across Reynolds numbers 2–6M and angles of attack from $5^{\circ}$ to $15^{\circ}$. ShapeNet-Car (chang2015shapenet) is a collection of 3D car geometries used for aerodynamic prediction, containing approximately 3700 mesh points per sample. DrivAerNet and DrivAerNet++ (elrefaie2024drivaernet) are distinct high-fidelity CFD datasets of parametric car geometries: DrivAerNet comprises 4,000 designs with approximately 500K surface vertices per car, while DrivAerNet++ extends this to over 10,000 designs with greater geometric diversity across multiple vehicle classes (SUVs, sedans, hatchbacks). All datasets model cars or airfoils as graphs with nodes representing surface mesh vertices and edges following mesh connectivity.
On the intermediate-scale benchmarks (AirfRANS and ShapeNet-Car, Table 4), GIST outperforms Transolver, achieving 0.0028 MSE on AirfRANS (vs. 0.0142) and 4.68 MSE on ShapeNet-Car (vs. 4.99). On the large-scale benchmarks (DrivAerNet and DrivAerNet++, Table 5), GIST achieves state-of-the-art on both: 20.10% relative $\ell_{2}$ error on DrivAerNet (vs. 20.35% for the previous best) and 18.60% on DrivAerNet++ (vs. 20.05%). GIST’s linear scaling enables direct processing of these 500K-node meshes without downsampling, while prior methods require lossy projection to regular grids or lower-dimensional latent spaces. GIST’s spectral attention provides global receptive fields in a single layer, which aids performance on all these tasks.
Table 4: Surface pressure prediction accuracy on AirfRANS and ShapeNet-Car benchmarks. Baseline is Transolver (wu2024transolver). MSE is reported directly. Lower is better.
Model AirfRANS (MSE $\downarrow$) ShapeNet-Car (MSE $\downarrow$)
Transolver 0.0142 4.99
GIST (ours) 0.0028 4.68
@@ 5 Conclusions
We presented GIST, a gauge-invariant spectral transformer whose contributions span three levels. First, we identified gauge invariance breaking as a fundamental failure mode shared by both exact and approximate spectral methods, causing catastrophic generalization loss in inductive settings. Second, we proved formally that gauge invariance is a necessary condition for discretization-invariant neural operators, establishing the first discretization mismatch bound $\mathcal{O}(n^{-1/(m+4)})+\mathcal{O}(r^{-1/2})$ for a graph neural operator. Third, we proposed an architecture that achieves gauge invariance, end-to-end $\mathcal{O}(N)$ complexity, and global spectral expressivity simultaneously, three properties that no prior method combines. By restricting attention to inner products of spectral embeddings, which are unbiased estimators of an exactly gauge-invariant kernel, GIST recovers the symmetry algorithmically while maintaining linear scaling.
@@
Appendix A Proofs
@@ A.1 Proof of Proposition Proposition: Gauge-Invariant Spectral Self-Attention
We prove that the Gauge-Invariant Spectral Self-Attention mechanism is discretization-invariant with quantifiable discretization mismatch error through three stages of analysis. The proof establishes that the attention kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ estimated by the spectral embeddings converges to a well-defined continuum kernel as mesh resolution increases, allowing us to bound the discretization mismatch error between any two discretizations of the same manifold.
@@
Gauge-Invariant Spectral Self-Attention is a discretization-invariant Neural Operator with bounded discretization mismatch error. Let $M$ be a compact $m$-dimensional Riemannian manifold and $\mathcal{G}_{n}$ a sequence of graphs obtained by sampling $n$ nodes from $M$ with $n\to\infty$. Let $f$ be a spectral filter applied to a graph operator $P_{n}$ of $\mathcal{G}_{n}$, and let $\phi_{i}^{n}=f(P_{n})_{i\cdot}^{\top}$ be the spectral embedding of node $i$. Then:
@@
    (Special case: Laplacian eigenmaps.) When $f(\lambda)=\lambda^{-1/2}$ and $P_{n}=\mathcal{L}_{n}$, the inner products $\langle\phi_{i}^{n},\phi_{j}^{n}\rangle$ converge to the Green’s function $G_{M}(x_{i},x_{j})$ of the Laplace-Beltrami operator at rate $\mathcal{O}(n^{-1/(m+4)})$. More generally, for any Lipschitz spectral filter $f$, the kernel $K_{n}(i,j)=\langle\phi_{i}^{n},\phi_{j}^{n}\rangle=[f(P_{n})f(P_{n})^{\top}]_{ij}$ converges to a well-defined continuum kernel $K_{\infty}(x_{i},x_{j})$ at the same rate (Corollary A.3).
@@
    (Discretization mismatch bound.) For any two discretizations $\mathcal{G}_{n}$ and $\mathcal{G}_{n^{\prime}}$ of the same manifold $M$ with $n\leq n^{\prime}$, the attention kernel values at corresponding points $x_{i},x_{j}\in M$ satisfy
    $\left|\langle\tilde{\phi}_{i}^{n},\tilde{\phi}_{j}^{n}\rangle-\langle\tilde{\phi}_{i}^{n^{\prime}},\tilde{\phi}_{j}^{n^{\prime}}\rangle\right|=\mathcal{O}\!\left(n^{-1/(m+4)}\right)+\mathcal{O}\!\left(r^{-1/2}\right),$
    where $\tilde{\phi}_{i}^{n}$ is the $i$-th row of $\widetilde{\Phi}^{n}=f(P_{n})\cdot R^{n}$. The two sources are independent: the first is the discretization mismatch error, the second the statistical estimation error from the random embedding.
  3. (iii)
    (Gauge invariance.) GIST’s attention kernel $K(i,j)=[f(P)f(P)^{\top}]_{ij}$ is exactly invariant under all gauge transformations of the eigenvector basis (sign flips, rotations within degenerate eigenspaces, solver choices). Combined with (ii), this ensures that learned parameters transfer across discretizations (discretization-invariance): the attention weights converge to a gauge-invariant continuum limit that does not depend on the arbitrary eigenvector conventions of any particular discretization.